Find the dimensions of a right triangle if the area is 120 square units and one leg is 4 more than twice the other leg.
Please help!!

Answer by edjones(8007)

a=2b+4
.5*a*b=A
.5(2b+4)*b=120
b(b+2)=120
b^2+2b-120=0 expand and subtract 120 from each side.
(b+12)(b-10)=0 Factor
b=10, a=24
b=-12 can't be an answer because it is negative.
Ed